# Information-Geometric Barycenters for Bayesian Federated Learning

### v2

@@ Abstract
Federated learning (FL) is a widely used and impactful distributed optimization framework that achieves consensus through averaging locally trained models. While effective, this approach may not align well with Bayesian inference, where the model space has the structure of a distribution space. Taking an information-geometric perspective, we reinterpret FL aggregation as the problem of finding the barycenter of local posteriors using a prespecified divergence metric, minimizing the average discrepancy across clients. This perspective provides a unifying framework that generalizes many existing methods and offers crisp insights into their theoretical underpinnings. We then propose BA-BFL, an algorithm that retains the convergence properties of Federated Averaging in non-convex settings. In non-independent and identically distributed scenarios, we conduct extensive comparisons with statistical aggregation techniques, showing that BA-BFL achieves performance comparable to state-of-the-art methods while offering a geometric interpretation of the aggregation phase. Additionally, we extend our analysis to Hybrid Bayesian Deep Learning, exploring the impact of Bayesian layers on uncertainty quantification and model calibration.
Index Terms:
Bayesian Federated Learning, Hybrid Bayesian DL, Uncertainty Quantification, Model Aggregation.
@@ I Introduction
Federated Learning (FL) has emerged as the de facto standard for decentralized learning, particularly in scenarios requiring strong privacy guarantees. As introduced in [1], an FL system consists of a central server that maintains a global model and interacts with multiple clients, i.e., end-user devices, each holding private local data. FL schemes generally operate in two phases: a local learning phase, where each client trains a local model using its own data, and an aggregation phase, where the locally updated models are sent to the server and combined according to a predetermined rule. These phases repeat iteratively, with the global model from the previous iteration serving as the updated local models in the current iteration.
@@
Aggregation plays a central role in FL, enabling the combination of individual client contributions into a global model while preserving data privacy and ensuring communication efficiency. Although various aggregation strategies have been proposed (e.g., [10]), most known aggregation methods rely on variants of weighted averaging. Notable examples include FedAvg [1] and FedProx [2], which aggregate local models by computing a weighted average of their parameters. The choice of weights is an important design parameter, encoding auxiliary attributes, such as the relative importance of each client’s model, for the overall objective or proportionally to the volume of data available at each client.
@@
A key challenge in FL, and generally in distributed learning systems, is the statistical heterogeneity among participating clients. In real-world scenarios, client datasets rarely satisfy the idealized assumption of independent and identically distributed (i.i.d.) data; instead, they exhibit varying degrees of heterogeneity or distributional shifts across clients. As surveyed in [11, 2], five major forms of heterogeneity are commonly identified: label distribution skew: variations in the representation of specific labels across clients (e.g., under- or over-represented classes); feature distribution skew: differences in the feature distributions associated with the same label; concept drift: the same labels correspond to different features across clients; concept shift: cases where identical samples are labeled differently by different clients, and quantity skew: differences in the number of data samples held by each client. While these heterogeneities are highly relevant in practice, addressing all of them simultaneously remains a major challenge. As a result, most existing FL research targets only a subset of these non-idealities [12, 13, 14]. Given the growing applicability of FL in real-world settings, uncertainty quantification and model calibration are central characteristics for building trustworthy and reliable models. Nonetheless, these aspects remain largely underexplored in the existing research around deterministic FL. A preliminary study on the topic, with specific application to healthcare, is presented in [15], providing an overview of various uncertainty quantification methods within the deterministic FL, later implemented in [16]. It is important to note, however, that the techniques discussed are mainly inspired by Bayesian approaches, including Bayesian ensembles and Monte Carlo dropout.
@@
Improving the reliability of models is where Bayesian learning excels, as Bayesian methods allow for better quantification of the model uncertainty and calibration, rendering it a compelling solution for FL in non-i.i.d. contexts. FedPPD [17] introduces an FL framework that incorporates uncertainty quantification, where each client, in each round, estimates the posterior distribution over its parameters and the posterior predictive distribution (PPD). The PPD is then distilled into a single deep neural network and sent to the server. pFedBayes [18] and Fedpop [19] are also Bayesian approaches with a focus on uncertainty quantification aspects, proposed in the context of personalized Bayesian Federated Learning (BFL). Nonetheless, we emphasize that many of the existing methods [18, 20, 17, 21] rely on variations of weighted averaging of the posteriors’ parameters.
@@
This work introduces a unifying perspective on aggregation methodologies within the context of BFL through the lens of barycentric aggregation (BA-BFL). Given a divergence metric, we interpret the aggregation process as a geometric problem, where the global model is identified as the barycenter of the local posteriors. In contrast to existing methods that often rely on heuristic variations of averaging in the parameter space, our approach is theoretically grounded, as it explicitly minimizes the average divergence between the global posterior and the local posteriors. We show that this methodology generalizes several aggregation strategies previously proposed in the literature. Furthermore, we show that the proposed methods preserve the convergence properties of FedAvg, even in non-convex settings (see Theorem III.1). We evaluate our techniques against state-of-the-art Bayesian aggregation methods in terms of accuracy and uncertainty quantification in a heterogeneous setting. The results demonstrate that our approach achieves performance comparable to existing statistical aggregation techniques. To bridge gaps in the BFL literature, and building on insights from Hybrid Bayesian Deep Learning (HBDL) [22], we extend our analysis to examine the impact of limiting the number of Bayesian layers on the performance of various Bayesian aggregation methods.
@@ II Background and Related Work
An FL system [1] consists of a central server and $N$ clients, engaging in an iterative learning process involving server-client communication. For each communication round, the $k^{th}$ client trains its local model, parameterized by $\theta_{k}$, on its private data $\mathcal{D}_{k}$. Subsequently, the model parameters $\theta_{k}$ are sent to the server, which aggregates all the clients’ models to obtain the global model, later sent to the clients to refine their local models for the next communication round. By aggregating the local models of the clients, FL seeks to achieve a global model $\theta^{*}$ on $\mathcal{D}=\bigcup_{k=1}^{N}\mathcal{D}_{k}$, the aggregated dataset from all participating clients.
In general, an FL system considers objective functions of the form
@@
where $f_{k}({\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}[\mathcal{L}(\theta;(x,y)]$ is the local objective function of the $k^{th}$ client and $w_{k}$ is its associated weight such that $\sum_{k=1}^{N}w_{k}=1$. Minimizing $f_{k}(\theta)$ at each communication round locally results in $\theta_{k}$.
@@
Despite the remarkable results in terms of model performance, deep learning does not address crucial problems in realistic scenarios, such as reliability and uncertainty quantification. In a recent position paper [24], the authors propose Bayesian Deep Learning as a solution for the ethical, privacy, and safety challenges of modern DL. Acknowledging ongoing issues with Bayesian DL, e.g., the added complexity cost of performing Bayesian learning on large-scale deep models, the authors envision the alternative framework of HBDL to maintain the efficiency and lower complexity of deep learning while retaining the reliability of Bayesian DL. HBDL is also discussed in [25] as Bayesian inference on the (n-)last layer(s) only. The idea behind this framework is to substitute part of the layers in a Bayesian deep model with deterministic layers, making them equivalent to the classical deep learning counterpart. The possibility of having only a partially Bayesian model allows for retaining uncertainty quantification capabilities while reducing the complexity compared to a fully Bayesian model.
@@
BFL aims to incorporate the strengths of Bayesian deep learning into the FL framework, offering a potential solution to the challenges outlined earlier. In this work, we focus primarily on parametric client-side BFL methods [26], and we categorize them according to their strategies for global model construction. Table II summarizes this discussion.
  • Multiplicative Aggregation of Posteriors: FedPA [6] employs Stochastic Gradient Markov Chain Monte Carlo (MCMC) for local posterior inference, aggregating client posteriors through a product of Gaussian distributions. FOLA [7] approximates local posteriors using Laplace approximation. A multivariate Gaussian product mechanism is used for global posterior construction, while prior distributions derived from the global posterior guide local training providing a continual learning setting. FedEP [8] frames FL as a distributed Variational Inference problem, aligning the global posterior with local posteriors using multiplicative aggregation. FedIvon [9] adopts Improved Variational Online Newton (IVON) [27] to approximate local posteriors as Gaussians with diagonal covariance, efficiently updating both mean and variance using second-order information. The server aggregates these local posteriors via the weighted product of the local Gaussians. It is important to note that approximating the global posterior as the weighted product of local posteriors is equivalent to computing their geometric mean, which also corresponds to the RKLB aggregation method introduced in this work, offering a geometric explainability to this aggregation’s strategy.
  • Weighted Averaging of Posteriors’ Statistics: pFedBayes [18] employs variational inference to incorporate uncertainty into model parameters. From a continual learning perspective, it minimizes the KL divergence between global and local posterior distributions, balancing local reconstruction error with global alignment, mainly presented in the paper as a personalization aspect. FVBA [20] investigates the aggregation of Variational Bayesian Neural Networks with five statistical aggregation schemes. In [21], the authors integrate Bayesian deep learning with FL, employing variational inference and Monte Carlo Dropout for local models’ inference. As in [20], the different statistical aggregations are evaluated, highlighting the importance of the choice of the aggregation method on the model performance. FedPPD [17] leverages MCMC sampling for local posterior inference and distilling posterior predictive distributions into single deep neural networks using Stochastic Gradient Langevin Dynamics. It adopts either simple averaging or a distillation-based approach for global aggregation.
  • Bayesian Posterior Inference: FedHB [23] introduces a hierarchical Bayesian framework where local model parameters are governed by a global latent variable. variational inference is used to optimize local and global posteriors through a block-coordinate optimization.
@@ III Proposed Method
[…]
@@ IV Experiments
We assign the aggregation weights to reflect the importance of each client in proportion to the volume of data locally owned, i.e., $w_{k}=\tfrac{|\mathcal{D}_{k}|}{|\mathcal{D}|}$ where ||$\cdot$|| ⋅ | indicates the number of samples in the dataset.
@@
Table 1 summarizes the accuracy performance of BA-BFL using RKLB and WB, alongside the considered baseline methods. All aggregation methods are evaluated under the same model architecture, with an equal number of Bayesian layers. The results show that the proposed aggregation methods achieve performance comparable to the baselines in most scenarios. Furthermore, all Bayesian methods consistently outperform FedAvg across all evaluated datasets. Notably, the best accuracy is obtained with a single Bayesian layer for FashionMINST and SVHN, and with two Bayesian layers for CIFAR-10. Interestingly, increasing the number of Bayesian layers does not always result in improved accuracy.
@@ V Conclusions and Future Work
In this paper, we introduced BA-BFL, a novel geometric interpretation of barycenters as a solution to the BFL aggregation problem, providing an explainable aggregation method. The information geometric interpretation of the aggregation step naturally enables operations such as clustering local posteriors directly on the statistical manifold, with potential applications in hierarchical FL. Building on this concept, we recovered two aggregation techniques based on the analytical results for Gaussian barycenters derived using two widely-used divergences: the squared Wasserstein-2 distance and the reverse KL divergence. We demonstrated that BA-BFL retains the same convergence properties as FedAvg for non-convex loss functions, performing robustly across both homogeneous and heterogeneous data scenarios. We experimentally evaluated the proposed methods in a heterogeneous setting, showing gains compared to state-of-the-art methods. Additionally, we examined the impact of varying the number of Bayesian layers in an HBDL context, assessing their effects on accuracy, uncertainty quantification, model calibration, and cost-effectiveness. For future work, we envision several extensions to our study, including expanding the family of distributions to incorporate non-parametric distributions and investigating alternative divergence measures. Furthermore, we plan to address the personalization problem within the framework of barycentric aggregation in BFL.

### v3

@@ Abstract
Federated learning (FL) is a widely used and impactful distributed optimization framework that achieves consensus by averaging locally trained models. While effective, this approach may not align well with Bayesian inference, where the model space is more naturally represented as a distribution space. Taking an information-geometric perspective, we reinterpret FL aggregation as the problem of finding the barycenter of local posteriors using a predefined divergence metric, minimizing the average discrepancy across clients. This perspective provides a unifying framework that generalizes many existing methods and offers crisp insights into their theoretical underpinnings. We then propose BA-BFL, an algorithm that retains the convergence properties of Federated Averaging in non-convex settings. In non-independent and identically distributed scenarios, we conduct extensive comparisons with statistical aggregation techniques, showing that BA-BFL achieves performance comparable to state-of-the-art methods while also providing a geometric interpretation of the aggregation phase. Additionally, we extend our analysis to Hybrid Bayesian Deep Learning, exploring the impact of Bayesian layers on uncertainty quantification and model calibration.
@@ I Introduction
Federated Learning (FL) has emerged as the de facto standard for decentralized learning, particularly in scenarios that demand strong privacy guarantees. As originally introduced in [1], an FL system consists of a central server that maintains a global model and interacts with multiple clients (end-user devices), each of which holds private local data. FL schemes typically operate in two phases. In the local learning phase, each client trains a model on its own private data. In the aggregation phase, the locally updated models are transmitted to the server and merged according to a predetermined rule. These phases repeat iteratively, with the global model from the previous iteration distributed to clients as the starting point for their local models in the current iteration.
@@
Aggregation plays a central role in FL, allowing individual client contributions to be combined into a global model while preserving data privacy and ensuring communication efficiency. Although various aggregation strategies have been proposed (e.g., [10]), most aggregation methods rely on variants of weighted averaging. Notable examples include FedAvg [1] and FedProx [2], which aggregate local models by computing a weighted average of their parameters. The choice of weights is an important design parameter, as it can encode auxiliary attributes such as the relative importance of each client’s model to the overall objective or reflect the amount of data available to each client.
@@
A key challenge in FL, and generally in distributed learning systems, is the statistical heterogeneity among participating clients. In real-world scenarios, client datasets rarely satisfy the idealized assumption of independent and identically distributed (i.i.d.) data. Instead, they often exhibit significant heterogeneity and distributional shifts across clients. As reviewed in [11, 2], five major forms of heterogeneity are typically identified: label distribution skew: differences in the frequency of specific labels across clients (e.g., under- or over-represented classes); feature distribution skew: differences in the feature distributions associated with the same label; concept drift: cases where the same label is associated with different feature distributions across clients; concept shift: cases where identical samples receive different labels from different clients, and quantity skew: differences in the number of data samples held by each client. While these heterogeneities are of high practical relevance, addressing all of them simultaneously remains a significant challenge. As a result, most of the existing FL research is focused on only a subset of these challenges [12, 13, 14]. Given the growing applicability of FL in real-world settings, uncertainty quantification and model calibration are central to building trustworthy and reliable models. Nonetheless, these aspects remain largely underexplored in existing research on deterministic FL. A preliminary study on the topic, with a specific application to healthcare, is presented in [15]. It provides an overview of various uncertainty quantification methods for deterministic FL, which were later implemented in [16]. However, it is important to note that the techniques discussed are primarily inspired by Bayesian approaches, such as Bayesian ensembles and Monte Carlo dropout.
@@
Bayesian learning excels at improving model reliability, as Bayesian methods enable more accurate uncertainty quantification and calibration, making them a compelling solution for FL in non-i.i.d. contexts. FedPPD [17] introduces an FL framework with built-in uncertainty quantification: in each round, each client estimates both the posterior distribution over its parameters and the posterior predictive distribution (PPD). The PPD is subsequently distilled into a single deep neural network, which is then sent to the server. pFedBayes [18] and Fedpop [19] are also Bayesian approaches with a focus on uncertainty quantification aspects, proposed in the context of personalized Bayesian Federated Learning (BFL). Nonetheless, we emphasize that many of the existing methods [18, 20, 17, 21] rely on variations of weighted averaging of the posteriors’ parameters.
@@
This work introduces a unifying perspective on aggregation methods in BFL through the lens of barycentric aggregation (BA-BFL). Given a divergence metric, we interpret the aggregation process as a geometric problem, where the global model is identified as the barycenter of the local posteriors. Unlike existing methods that often rely on heuristic variations of parameter averaging, our approach is theoretically grounded: it explicitly minimizes the average divergence between the global posterior and the local posteriors. We show that this methodology generalizes several aggregation strategies previously proposed in the literature. Furthermore, the proposed methods preserve the convergence properties of FedAvg, even in non-convex settings (see Theorem III.1). We evaluate our approach against state-of-the-art Bayesian aggregation methods, comparing accuracy and uncertainty quantification in heterogeneous settings. The results demonstrate that our method achieves performance comparable to existing statistical aggregation techniques. To bridge gaps in the BFL literature and buildinsights from Hybrid Bayesian Deep Learning (HBDL) [22], we further examine how limiting the number of Bayesian layers affects the performance of different Bayesian aggregation methods.
@@ II Background and Related Work
An FL system [1] consists of a central server and $N$ clients that engage in an iterative learning process through server-client communication. For each communication round, the $k^{th}$ client trains its local model, parameterized by $\theta_{k}$, on its private data $\mathcal{D}_{k}$. Subsequently, the model parameters $\theta_{k}$ are sent to the server, which aggregates them to obtain the global model. The updated global model is then distributed back to the clients to refine their local models in the next communication round. Through this process, FL aims to learn a global model $\theta^{*}$ on the aggregated dataset $\mathcal{D}=\bigcup_{k=1}^{N}\mathcal{D}_{k}$ from all participating clients.
In general, the objective function in FL takes the form
@@
where $f_{k}({\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}[\mathcal{L}(\theta;(x,y))]$ is the local objective function of the $k^{th}$ client, and $w_{k}$ is its associated weight, with $\sum_{k=1}^{N}w_{k}=1$. At each communication round, minimizing $f_{k}(\theta)$ locally produces the client update $\theta_{k}$.
@@
Despite its remarkable performance, deep learning does not address crucial challenges in realistic scenarios, such as reliability and uncertainty quantification. In a recent position paper [24], the authors propose Bayesian Deep Learning as a solution to the ethical, privacy, and safety challenges of modern deep learning. Acknowledging ongoing challenges with Bayesian DL, such as the additional computational cost of applying Bayesian methods to large-scale deep models, the authors envision the alternative framework of HBDL to preserve the efficiency and lower complexity of deep learning while retaining the reliability of Bayesian DL. HBDL is also discussed in [25] as Bayesian inference applied only to the last (or last few) layers. The core idea is to replace some of the layers in a Bayesian deep model with deterministic ones, thereby making the model closer to its classical deep learning counterpart. This partial Bayesian formulation makes it possible to retain uncertainty quantification capabilities while reducing complexity relative to a fully Bayesian model.
@@
BFL aims to incorporate the strengths of Bayesian deep learning into the FL framework and to provide a potential solution to the challenges outlined above. In this work, we focus primarily on parametric, client-side BFL methods [26]. We categorize these methods based on their strategies for global model construction. Table II provides a summary of this discussion.
  • Multiplicative Aggregation of Posteriors: FedPA [6] employs Stochastic Gradient Markov Chain Monte Carlo (MCMC) for local posterior inference, aggregating client posteriors through a product of Gaussian distributions. FOLA [7] approximates local posteriors using Laplace approximation. A multivariate Gaussian product mechanism is used for global posterior construction, while prior distributions derived from the global posterior guide local training, thereby enabling a continual learning setting. FedEP [8] frames FL as a distributed variational inference problem, aligning the global posterior with local posteriors through multiplicative aggregation. FedIvon [9] adopts Improved Variational Online Newton (IVON) [27] to approximate local posteriors as Gaussians with diagonal covariance, efficiently updating both mean and variance with second-order information. The server aggregates these local posteriors via the weighted product of the local Gaussians. It is important to note that approximating the global posterior as the weighted product of local posteriors is equivalent to computing their geometric mean, which also corresponds to the RKLB aggregation method introduced in this work, offering a geometric interpretation of this aggregation strategy.
  • Weighted Averaging of Posteriors’ Statistics: pFedBayes [18] employs variational inference to incorporate uncertainty into model parameters. From a continual learning perspective, it minimizes the KL divergence between global and local posterior distributions, balancing local reconstruction error with global alignment, which the paper presents mainly as a personalization aspect. FVBA [20] investigates aggregating variational Bayesian neural networks using five statistical aggregation schemes. In [21], the authors integrate Bayesian deep learning with FL, employing variational inference and Monte Carlo Dropout for inference in local models. As in [20], different statistical aggregations are evaluated, highlighting the importance of the aggregation method chosen for model performance. FedPPD [17] leverages MCMC sampling for local posterior inference and distills posterior predictive distributions into individual deep neural networks via Stochastic Gradient Langevin Dynamics. It adopts either simple averaging or a distillation-based global aggregation approach.
  • Bayesian Posterior Inference: FedHB [23] introduces a hierarchical Bayesian framework in which local model parameters are governed by a global latent variable. Variational inference is used to optimize the local and global posteriors through block-coordinate optimization.
@@ III Proposed Method
$p^{*}_{D_{\alpha}}=\frac{\left(\sum_{k=1}^{N}w_{k}p_{k}^{\alpha}\right)^{\frac{1}{\alpha}}}{\int\left(\sum_{k=1}^{N}w_{k}p_{k}^{\alpha}\right)^{\frac{1}{\alpha}}d\nu}.$ (5)
@@ IV Experiments
We assign the aggregation weights to reflect the importance of each client in proportion to the volume of data locally owned, i.e., $w_{k}=\tfrac{|\mathcal{D}_{k}|}{|\mathcal{D}|}$ where $|\cdot|$ indicates the number of samples in the dataset.
@@
Table 1 summarizes the accuracy performance of BA-BFL using RKLB and WB, alongside the considered baseline methods. All aggregation methods are evaluated under the same model architecture, with an equal number of Bayesian layers. The results show that the proposed aggregation methods achieve performance comparable to the baselines in most scenarios. Furthermore, all Bayesian methods consistently outperform FedAvg across all evaluated datasets. Notably, the best accuracy is obtained with a single Bayesian layer for FashionMNIST and SVHN, and with two Bayesian layers for CIFAR-10. Interestingly, increasing the number of Bayesian layers does not always result in improved accuracy.
@@ V Conclusions and Future Work
In this paper, we introduced BA-BFL, a novel geometric interpretation of barycenters as a solution to the BFL aggregation problem. This approach provides an explainable aggregation method. The information-geometric view of the aggregation step naturally enables operations such as clustering local posteriors directly on the statistical manifold, which has potential applications in hierarchical FL. Building on this concept, we recovered two aggregation techniques based on analytical results for Gaussian barycenters using two widely used divergences: the squared Wasserstein-2 distance and the reverse KL divergence. We demonstrated that BA-BFL retains the convergence properties of FedAvg for non-convex loss functions and performs robustly in both homogeneous and heterogeneous data scenarios. We experimentally evaluated the proposed methods in heterogeneous settings, showing improvements over state-of-the-art methods. We also examined the impact of varying the number of Bayesian layers in an HBDL context, evaluating their effects on accuracy, uncertainty quantification, model calibration, and cost-effectiveness. For future work, we envision several extensions, including expanding the family of distributions to include non-parametric ones and exploring alternative divergence measures. We also plan to address the personalization problem within the barycentric aggregation framework for BFL.
Acknowledgments
The work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation programme (Grant agreement No.101003431), the IMT ”Futur, Ruptures & Impacts” program, and the European Commission through the Horizon Europe/JU SNS project, ROBUST-6G (Grant Agreement No. 101139068).